suppose that in given year in closed economy with goverment,private and public sector enterpries produced goods and services worth 10 lakhs i.e aggregate GVA in the given year was rs 10 lakhs . they sold in th same year goods and service of rs 8 lakhs to households and rs 5 lakh to goverment administration and defence . besides goverment administration and defence paid rs 10 lakhs in wages and salaries and rs 1 lakh in interest to households. what is the aggregate gross investment of the firms in the given year? calculate GDP of the economy by the expenditure method.


1
Expert's answer
2021-06-10T10:59:45-0400

Solution:

1.). The aggregate gross investment of the firms in the given year refers to the total investor spending on business capital goods.

The aggregate gross investment = 5 lakhs

 

2.). GDP using the expenditure method (Y) = C + I + G + X – M

Consumption (C) = 8 + 1 = 9

Government spending (G) = 10 = 10

Aggregate gross investment (I) = 5

GDP using the expenditure method (Y) = 9 + 10 + 5 = 24

 

GDP using the expenditure method (Y) = 24 lakhs